OpenStudy (anonymous):

The slope of a line is 2, and the y-intercept is 0. What is the equation of the line written in slope-intercept form? y = x + 2 y = 2 y = 2x C

OpenStudy (jdoe0001):

Twis7ed remember that slope intercept form means y = mx + b where m is the slope of the line and b is the y-intercept. Since you know m = 0 and b = 6 you can just plug these into y = mx + b and simplify. --------------------- do the same

OpenStudy (anonymous):

so c like i said

OpenStudy (jdoe0001):

\(\begin{matrix} y = &\boldsymbol{ m}&x &+\boldsymbol{ b}\\ & slope && y-intercept \end{matrix}\)

OpenStudy (anonymous):

so c like i said

OpenStudy (jdoe0001):

let's see it, how did you get C?

OpenStudy (anonymous):

slope is 2 y-i=0 therefore y=2x

OpenStudy (jdoe0001):

\(\begin{matrix} y = &m&x &+ b\\ & slope && y-intercept\\ y =& 2&x&+0 \end{matrix}\\ \implies \color{blue}{y = 2x}\)

OpenStudy (jdoe0001):

yes, you're correct :)

OpenStudy (anonymous):

like i sadi eight hours ago C....thanks for helping

OpenStudy (anonymous):

said*

OpenStudy (dan815):

GENIUS

OpenStudy (dan815):

you have a very bright future

OpenStudy (dan815):

:D